The perimeter of a triangle is 153 ft. Find the length on each side:
(4x) ft on left side of triangle
(6x+1) ft on right side

(6x+8) ft bottom of triangle.

Answer by stanbon(75887)

The perimeter of a triangle is 153 ft. Find the length on each side:
(4x) ft on left side of triangle
(6x+1) ft on right side
(6x+8) ft bottom of triangle.
------
Equation:
4x + 6x+1 + 6x+8 = 153
16x + 9 = 153
----
16x = 144
x = 9
-----
left = 4x = 4*9 = 36 ft
----
right = 6x+1 = 6*9 + 1 = 55 ft
-----
bottom = 6x+8 = 6*9 + 8 = 62 ft
===================================
Cheers,
Stan H.
---------------------